OpenStudy (anonymous):

Find the integral integral of -sec^2 x dx

zepdrix (zepdrix):

Super secrit hint!! :O Do you remember this derivative?\[\Large\rm \frac{d}{dx}\tan x=?\]

OpenStudy (anonymous):

yes, sec^2 is the derivative

zepdrix (zepdrix):

Good good good. So the anti-derivative of sec^2x is tanx, yes? And just carry the negative along for the ride.

OpenStudy (anonymous):

but I got -tanx + C as an answer and got it wrong :( ?

zepdrix (zepdrix):

Mmmmm you shouldn't have gotten it wrong... The read the question correctly? Or is it a definite integral with bounds?

zepdrix (zepdrix):

You read*

OpenStudy (anonymous):

It just says find the integral, I found the indefinite one for this, do you think it can be done for definite integral?

zepdrix (zepdrix):

Well, ya if you're given bounds it can. Like you might be given limits of pi/4 to 3pi/4 or something like that in radians. But if you weren't given anything like that, it's fine. Maybe just a formatting error? Do you need to put the x in brackets or something? -tan(x)+C

OpenStudy (anonymous):

nope, that is just what I got, let me check again.

OpenStudy (anonymous):

No, that's it... maybe its some error in the exercise :/ that's ok, thanks anyways.

zepdrix (zepdrix):

weird :U

OpenStudy (anonymous):

This is how I would go about it. To solve \[\frac {d}{dx}\sec^2x\] I would use the chain rule where \(g(x)=\sec x\) and \(f(x)=x^2\) \[\frac{d}{dx}f(g(x))=g'(x)f'(g(x))\] Popping in \(g(x)\) and \(f(x)\) gives us \[\frac{d}{dx}f(\sec x ))=(\sec x \tan x )f'(\sec x)\] \[\frac{d}{dx} \sec^2 x=(\sec x \tan x)(2 \tan x)\] \[\frac{d}{dx} \sec^2 x= 2 \tan^2 x \sec x\]

OpenStudy (anonymous):

Ok, i'll do both ways and see which one would be correct. Thank you!

zepdrix (zepdrix):

Hmm that's the derivative, not the integral Saleh. Maybe that's what the question wanted though heh.

OpenStudy (anonymous):

I think that's my eyes telling me to get some sleep!

OpenStudy (anonymous):

To sum up what zepdrix mentioned, \[\frac{d}{dx} \tan x = sec^2x\] Or \[d \tan x = \sec^2 x dx\] Then take a anti-derivative (or integral) from both sides \[\int d \tan x = \int \sec^2 x dx\] \[\tan x + c= \int \sec^2 x dx\] In your case \[-\int \sec^2 x dx = -\tan x + c \]